### v2

@@ 2.2 Uncertainty from a generalized bias-variance perspective
$+\underbrace{\mathbb{E}_{\mathcal{D}_{N},\gamma}\big[D_{\ell}(\bar{p}^{*}(y|\boldsymbol{x}),\hat{p}_{N,\gamma}(y|\boldsymbol{x})\big]}_{\text{generalized variance}},$ (3)
where $\bar{p}^{*}(y|\boldsymbol{x})=\operatorname*{\arg\min}_{p\in\mathbb{P}(\mathcal{Y})}D_{\ell}(p,\hat{p}_{N,\gamma}(y|\boldsymbol{x}))$ the dual mean of $\hat{p}_{N,\gamma}(y|\boldsymbol{x})$ under data- and procedural uncertainty. The first term in Eqn. 2.2 corresponds to the intrinsic variability within the data-generating process, and hence to aleatoric uncertainty. The last term reflects the generalized variance of the predictive distribution under data and procedural uncertainty, and ca,n under suitable conditions of the loss function $\ell$ [31] be further decomposed by means of the conditional bias-variance decomposition for Bregman divergences [30] as
@@
$\mathbb{E}_{\mathcal{D}_{N},\gamma}\big[D_{\ell}(\bar{p}^{*}(y|\boldsymbol{x}),\hat{p}_{N,\gamma}(y|\boldsymbol{x})\big]$ $=\underbrace{\mathbb{E}_{\mathcal{D}_{N}}\big[D_{\ell}(\bar{p}^{*}(y|\boldsymbol{x}),\bar{p}_{|\mathcal{D}_{N}}^{*}(y|\boldsymbol{x}))\big]}_{\text{procedural uncertainty}}$
$+\underbrace{\mathbb{E}_{\mathcal{D}_{N},\gamma}\big[D_{\ell}(\bar{p}_{|\mathcal{D}_{N}}^{*}(y|\boldsymbol{x}),\hat{p}_{N,\gamma}(y|\boldsymbol{x}))\big]}_{\text{data uncertainty}}.$ (4)
Finally, the second term in Eqn. 2.2 reflects the generalised bias, which is the difference between the mean predicted distribution $\bar{p}^{*}(y|\boldsymbol{x})$ under data and model uncertainty and the true conditional distribution $p(y|\boldsymbol{x})$. For universal approximators, such as many types of deep neural networks, the model uncertainty is zero [32, 33, 34]. Such models are asymptotically unbiased, but in the finite data regime, a bias is unavoidable to make learning possible [20]. For more restrictive hypothesis spaces, the bias even does not disappear when the training dataset becomes infinitely large. Thus, in both cases, the bias term cannot be ignored when analysing epistemic uncertainty. We see that any proper scoring rule whose associated divergence can be related to a Bregman divergence allows for the split of its expected risk into an aleatoric and an epistemic part, where the epistemic part can be further split into a part representing bias, procedural and data uncertainty, respectively. In the next sections we will look at this decomposition more closely for the regression case, using mean squared error and negative log likelihood as loss functions.
3 Evaluating epistemic uncertainty via statistical simulations
3.1 The reference distribution
Quantifying epistemic uncertainty objectively is challenging, and evaluating the correctness of such quantifications is equally difficult. While calibration tests exist for aleatoric uncertainty in probabilistic models [35], epistemic uncertainty lacks established evaluation protocols and is often assessed indirectly through downstream tasks like out-of-distribution detection [5, 22, 36], robustness to adversarial attacks [37, 38], and active learning [39]. Correctness of second-order epistemic uncertainty representations can only be evaluated through statistical simulations, in which bias and variance can be analysed individually. To this end, Jürgens et al. [24] recently introduced the notion of a reference distribution, which is the theoretical distribution over first-order models if training data could be resampled. Here we further extend this notion by taking the variance due to procedural uncertainty into account:
Definition 1.
@@ 3.1 The reference distribution
$q_{N,\gamma}(\boldsymbol{\theta}|\boldsymbol{x},\phi):=\mathbb{P}(\boldsymbol{\hat{\theta}}_{N,\gamma}(x)=\boldsymbol{\theta})=\int_{\Gamma}\int_{\mathcal{X}\times\mathcal{Y}}\mathbb{I}\Big[(\boldsymbol{\hat{\theta}}_{N,\gamma}(x)=\boldsymbol{\theta})\Big]dP^{N}dP^{\gamma},$ (5)
@@
Intuitively, the reference distribution describes the uncertainty in the first-order parameters, induced by having a limited dataset and procedural uncertainty due to different weight initializations. In simulation studies, one can estimate the reference distribution by drawing $n_{d}$ training datasets $\mathcal{D}_{N_{1}},\dots,\mathcal{D}_{N_{n_{d}}}$ and $n_{\gamma}$ procedural parameters $\gamma_{1},\dots,\gamma_{n_{\gamma}}$ from $P^{N}$ and $P^{\gamma}$, respectively, and calculating the empirical distribution function. The (generalized) variance of the reference distribution describes the variability due to limited data expressivity as well as the procedural variance, and can be decomposed by means of the decomposition in Eqn. 2.2. In addition, in statistical simulations, where the data-generating process $p(y|\boldsymbol{x})$ is known, the bias of the reference distribution can also be measured.
3.2 Bias-variance decomposition in the regression case
We start this section by deriving expressions for aleatoric and epistemic uncertainty estimates according to Section 2.2 for the regression case, which is the setting we focus on for the remainder of the paper. Regression has the advantage over classification that our key findings can be more easily and intuitively visualized, and has also not been studied as extensively as the classification case for uncertainty disentanglement.
We consider the following data-generating process: $y_{i}=f(\boldsymbol{x}_{i})+\epsilon_{i}$, where $\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2}(\boldsymbol{x}_{i}))$ is data-dependent (heteroscedastic) noise, and $f$ is a deterministic function, such that $p(\cdot|\boldsymbol{x})$ forms the density of a Gaussian distribution with mean $f(\boldsymbol{x})$ and variance $\sigma^{2}(\boldsymbol{x})$. For the case where the loss function is the mean squared error loss, $\ell_{\text{MSE}}(\hat{y},y)=(\hat{y}-y)^{2}$, Eqn. 2.2 then becomes the classical bias-variance decomposition:
$\mathbb{E}_{y|\boldsymbol{x}}\big[\mathbb{E}_{\mathcal{D}_{N},\gamma}(\hat{y}-y)^{2}|\boldsymbol{x}\big]=\sigma^{2}(\boldsymbol{x})+\text{Var}_{\mathcal{D}_{N},\gamma}(\hat{y}|\boldsymbol{x})+\text{bias}^{2}\big[\hat{y}|\boldsymbol{x}\big],$ (6)
where the variance of the prediction $\hat{y}$ can be further split into procedural and data uncertainty by means of the law of total variance:
$\text{Var}_{\mathcal{D}_{N},\gamma}(\hat{y}|\boldsymbol{x})=\mathbb{E}_{\mathcal{D}_{N}}\big[\text{Var}_{\gamma}(\hat{y}|\boldsymbol{x})|\mathcal{D}_{N}\big]+\text{Var}_{\mathcal{D}_{N}}(\mathbb{E}_{\gamma}[\hat{y}|\boldsymbol{x}]|\mathcal{D}_{N}).$ (7)
3.3 Simulating data and procedural uncertainty
Figure 2: Estimated procedural and data uncertainty of the reference distribution for increasing sample sizes. The reference distribution is estimated by resampling training data and procedural parameters, using $n_{\gamma}=10$ and $n_{d}=20$. Procedural and data uncertainty are calculated using the decomposition formula in Eqn. 2.2. Even when the model predictions (complete line) have an important deviation (bias) from the ground-truth expected value (dashed line), the epistemic uncertainty estimate decreases with more data.
In the following we illustrate how one can use the reference distribution in Def. 1 to calculate expressions for data- and procedural uncertainty in the regression case. To this end, we adapt the synthetic regression setting from Rossellini et al. [40]. Precisely, we generate datasets $\mathcal{D}_{N}=\{(x_{i},y_{i})\}_{i=1}^{N}$, where
$x_{i}\sim Beta(1.2,0.5),\quad y_{i}=\sin(\frac{1}{5(x_{i}+0.16)^{3}})+\epsilon_{i},\quad\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2}(x_{i})),$ (8)
with $\sigma^{2}(x)=x^{4}$. The setup can be seen in Figure 1(a). This dataset highlights the differences between aleatoric and epistemic uncertainty: for low $x$ values, the complexity of the underlying data-generating function is high, while data is limited, leading to high epistemic uncertainty. On the other hand, for $x$ close to $1$, aleatoric uncertainty is high, while simultaneously having more data points and a less complex underlying function, implying low epistemic uncertainty. In order to estimate the reference distribution we train models with $n_{d}$ training datasets and $n_{\gamma}$ random weight initializations. As the underlying model class for obtaining the individual risk minimizers $\hat{p}_{\gamma,N}$, we use an MLP. Details about the experimental setup are discussed in Appendix A.1111The implementation for all our examples can be found in this GitHub repository.. Individual estimates for data- and procedural uncertainty can be obtained by the decomposition formula in Eqn. 2.2. The bias can be calculated as the absolute distance between the mean prediction and the underlying ground truth conditional expectation.
Figure 2 shows the obtained procedural and data uncertainty of the reference distribution for different training dataset sizes. First, we see that the total variance of the reference distribution decreases when sample size increases, supporting the claim that the reference distribution properly represents data uncertainty. Second, we see that the estimate of the procedural uncertainty also decreases with an increasing sample size, which is not unexpected, because initial procedural configurations become less important when more data is observed. Third, in the region with high epistemic uncertainty, the reference distribution has a very large bias, due to the underlying complexity of the data generating function. The bias is in fact so high that aleatoric uncertainty cannot be correctly estimated anymore. This last finding confirms our central position: Any claim to properly represent epistemic uncertainty must examine not only how different sources of uncertainty are modeled, but also whether bias is systematically distorting the partitioning of uncertainty. When the bias is high, estimates of both epistemic and aleatoric uncertainty can be completely misleading.
4 Which part of epistemic uncertainty do current methods capture?
Throughout this section we evaluate methods for the heteroscedastic regression setting using the classical variance-based split of the predictive variance as in Eqn. 7. We hereby focus on the mean squared error case of the more general pointwise decomposition in Section 2.2, which is the most commonly used decomposition for regression in practice. Departing from a Bayesian perspective, we first describe in this section how various second-order methods claim to obtain disentangled estimates for aleatoric and epistemic uncertainty. Secondly, we analyse two representative methods that approximate the posterior distribution over the parameters in more detail, illustrating that they tackle (and neglect) different sources of epistemic uncertainty. In the last part of this section as well as in Appendix A.1, we present experimental results for eight selected methods, highlighting that all of them have difficulties to model the different sources of epistemic uncertainty in a complete manner.

### v3

@@ 3 Epistemic uncertainty estimation methods are fundamentally incomplete
$+\underbrace{\mathbb{E}_{\mathcal{D}_{N},\gamma}[D_{\ell}(\bar{p}^{*}(y|x),\hat{p}_{N,\gamma}(y|\boldsymbol{x}))]}_{\text{generalized variance}}.$ (7)
The first term in Eqn. 3.1 represents the intrinsic noise of the data-generating process, and therefore corresponds to aleatoric uncertainty, which is irreducible with more training data. The second term captures the generalised bias, defined as the divergence between the mean predicted distribution $\bar{p}^{*}(y|\boldsymbol{x})$ averaged over data and procedural uncertainty, and the true conditional distribution $p(y|\boldsymbol{x})$. Finally, the third term reflects the generalized variance of the predictive distribution under data and procedural uncertainty. In Section 3.2, we will further decompose it into data and procedural components, by means of the conditional bias-variance decomposition for Bregman divergences (see (Gneiting and Raftery, 2007; Adlam et al., 2022)).
Connecting the bias component to Eqn. 5, one can decompose it into approximation error and estimation bias. Mainly the approximation error has seen some discussion in previous research on epistemic uncertainty, because the approximation error is nonzero when the ground-truth model is not in the hypothesis space, which might happen for certain hypothesis spaces, such as linear models. However, for universal function approximators, such as many types of deep neural networks, the approximation error will vanish (Hornik et al., 1989; Cybenko, 1989; Allenby and Labuschagne, 2009). Therefore, we believe that estimation bias deserves more discussion in modern deep learning research. It arises from the use of finite training datasets and from regularization procedures employed during training, which is inevitable to make learning possible with a finite dataset (Hastie et al., 2009). Approximation error and estimation bias are both not included in existing epistemic uncertainty estimators, such as the one in Eqn. 3, supporting our position that epistemic uncertainty estimation is fundamentally incomplete.
In the classical bias-variance decomposition for regression, the first term of Eqn. 3.1 is commonly estimated using Eqn. 3 as $\mathbb{E}[\hat{\sigma}^{2}]$, where $\hat{\sigma}^{2}$ denotes the predictive variance learned from the model. When minimizing squared loss, this quantity does not recover the true aleatoric noise (the details can be found in Appendix B.1). Instead, it satisfies $\mathbb{E}[\hat{\sigma}^{2}]=\mathbb{E}[(y-\hat{f}(x))^{2}]+\sigma^{2}$, as shown in (Zhang et al., 2024). Consequently, the learned aleatoric term absorbs observation noise and systematic prediction error: whenever $\hat{f}(x)$ deviates from the conditional mean, that deviation contributes to the learned $\hat{\sigma}^{2}$, inflating the aleatoric uncertainty estimation. Bias in the estimation of the mean will result in an overestimation of aleatoric noise.
Running Example - Deep Ensembles Figure 3 presents the uncertainty estimates produced by a Deep Ensemble on the synthetic dataset that we described earlier. The figure illustrates two key failure modes. First, in the top panel, the left side of the input domain exhibits a systematic prediction error: due to regularization, the model has not yet recovered the true function in this region. Despite the observation noise being close to zero, the estimated aleatoric uncertainty is inflated, indicating that bias in the mean predictor is being absorbed into the aleatoric term. Second, in the bottom panel, the epistemic uncertainty remains low in the same region, even though the predictor is clearly misspecified with respect to the true function. Taken together, these results show that, in the presence of systematic bias, deep ensembles can simultaneously overestimate aleatoric uncertainty and underestimate epistemic uncertainty, yielding an epistemic estimate that is inherently incomplete.
Figure 3: Deep Ensemble uncertainty estimates. The top row shows the training data points (blue crosses), the true conditional mean (dashed line), the model predictions (complete line), the $95\%$ confidence interval for aleatoric uncertainty (purple), as well as the true $95\%$ confidence interval of the aleatoric uncertainty (gray). The bottom row shows the estimate of epistemic uncertainty (green).
Despite a complex function on the left side of the domain, and a high aleatoric uncertainty on the right side, the synthetic example that we consider is with a one-dimensional feature space very simple. However, if uncertainty estimation is already unreliable in such a simple setting, there is very few hope that things get better for modern high-dimensional datasets. In Appendix C.3 we observe exactly the same issues for a real-world regression dataset. Moreover, for other second-order methods, estimation bias also leads to an overestimation of aleatoric uncertainty — see Appendix C.2 for additional experimental results that support this claim. In Bayesian methods, for instance, estimation bias is enforced by placing priors on the weights of the neural network. For simple hypotheis spaces, this allows for a flexible way of incorporating domain knowledge, but the choice of priors becomes quite arbitrary in large Bayesian neural networks (Wilson and Izmailov, 2020; Fortuin, 2022). Other methods, such as Monte Carlo Dropout and Bootstrap ensembles, also contain regularization techniques by design, which also lead to estimation bias.
3.2 Data and procedural components of epistemic uncertainty
In supervised learning, a trained predictor is a function of both the observed dataset and the learning procedure. A learning algorithm based on a neural network returns a predictive distribution $\hat{p}_{D_{N},\gamma}(y|\boldsymbol{x})$, which is dependent on both the dataset $D_{N}$ as well as the procedural parameters $\gamma$. As a result, variance can be decomposed into a procedural component, formed by the expected variance induced by $\gamma$, while the data is fixed, and a data component, consisting of the variance induced by resampling datasets $D_{N}$. Using the law of total variance, for $\hat{y}\sim\hat{p}_{\mathcal{D}_{N},\gamma}(y|\boldsymbol{x})$ we have
@@ 3.2 Data and procedural components of epistemic uncertainty
$\begin{split}\text{Var}_{\mathcal{D}_{N},\gamma}(\hat{y}|\boldsymbol{x})&=\mathbb{E}_{\mathcal{D}_{N}}\big[\text{Var}_{\gamma}(\hat{y}|\boldsymbol{x})|\mathcal{D}_{N}\big]\\ &\quad+\text{Var}_{\mathcal{D}_{N}}(\mathbb{E}_{\gamma}[\hat{y}|\boldsymbol{x}]|\mathcal{D}_{N}).\end{split}$ (8)
So far we claimed that second-order models ignore the bias component when estimating epistemic uncertainty. Using Deep Ensembles as a running example, we also argue that common second-order methods only track parts of the epistemic uncertainty that originates from variance.
Running Example - Deep Ensembles
Deep Ensembles capture procedural uncertainty only.
Let $\gamma\sim P^{\gamma}$ denote the random seed (weight initialization, data ordering, etc.) that differentiates each ensemble member’s training, where all members are trained on the same dataset $\mathcal{D}_{N}$. For the regression setting and a Deep Ensemble of size $d$, each member produces predictive parameters $\boldsymbol{\theta}{\gamma_{i}}(\boldsymbol{x})=(f_{\gamma_{i}}(\boldsymbol{x}),\sigma^{2}_{\gamma_{i}}(\boldsymbol{x}))$, for $i=1,\dots,d$. The empirical second-order distribution is a mixture of Dirac measures: $q(\boldsymbol{\theta}|x)=\frac{1}{d}\sum_{i=1}^{d}\delta_{\theta_{\gamma_{i}}(x)}(\theta).$ Applying the variance-based decomposition (Eqn. 3) and taking $d\to\infty$, the epistemic estimate is: $\text{Var}_{\gamma}\big(\mathbb{E}_{y|\boldsymbol{x}}[\hat{y}|\boldsymbol{x},\gamma]\big)=\mathbb{E}_{\gamma}\left[(f_{\gamma}(\boldsymbol{x})-\bar{f}(\boldsymbol{x}))^{2}\right],$ where $\bar{f}(\boldsymbol{x})=\mathbb{E}_{\gamma}[f_{\gamma}(\boldsymbol{x})]$. This term is linked to the first term in Eqn. 8, reflecting the procedural uncertainty. The data uncertainty component $\text{Var}_{D_{N}}(\mathbb{E}_{\gamma}[\hat{y}|\boldsymbol{x}]|\mathcal{D}_{N})$, which captures how predictions would vary under different training datasets from the same distribution, is entirely absent from the Deep Ensemble’s epistemic estimate.
While Deep Ensembles might explore multiple (local) optima of the training loss by using different random seeds (Wilson and Izmailov, 2020), Bayesian approaches attempt to explicitly approximate the posterior over parameters, which can yield more principled estimates of epistemic uncertainty. However, the practical methods used to approximate the posterior distribution also face restrictions: Approximate inference methods like the Laplace approximation or variational inference often capture only a single local mode (Wilson and Izmailov, 2020), employ simplifying assumptions (Psaros et al., 2023), or can be computationally costly in high-dimensional weight spaces. Other issues arise with methods that try to estimate the second-order distribution directly via loss minimisation, such as Deep Evidential Regression (Amini et al., 2020).
The theoretical analysis above raises an immediate question: Given a second-order uncertainty estimation method, how can one assess which components of epistemic uncertainty it represents? In the absence of observable ground-truth uncertainty, epistemic uncertainty is typically evaluated through downstream tasks (Gawlikowski et al., 2023): out-of-distribution detection (Hendrycks and Gimpel, 2017; Ovadia et al., 2019; Kopetzki et al., 2021), robustness to adversarial attacks (Malinin and Gales, 2019; Smith and Gal, 2018), selective prediction (Geifman and El-Yaniv, 2017) or active learning (Nguyen et al., 2022). These evaluation paradigms have important limitations. First, they conflate multiple uncertainty sources: out-of-distribution detection performance depends on both epistemic uncertainty quality and the specific distributional shift (Ovadia et al., 2019), while selective prediction benefits from total uncertainty rather than epistemic uncertainty alone (Hofman et al., 2025). Second, task performance does not isolate whether the decomposition is faithful, since a method may achieve good performance through mechanisms unrelated to true epistemic uncertainty. More fundamentally, Li et al. (2025) argue that out-of-distribution detection methods may be generally ill posed: uncertainty-based methods conflate high uncertainty with being far from training data, while feature-based methods conflate feature-space distance with it. Finally, none of these evaluation methods assess whether epistemic uncertainty captures the components we identify as missing.
We argue that separately quantifying procedural and data uncertainty is only possible through statistical simulation, where repeated sampling from the data-generating process can be performed or approximated. Following Jürgens et al. (2024), we formalize this through the notion of a reference distribution.
Definition 3.1.
@@
$q_{N,\gamma}(\boldsymbol{\theta}|\boldsymbol{x},\phi)$ $:=\mathbb{P}(\boldsymbol{\hat{\theta}}_{N,\gamma}(x)=\boldsymbol{\theta})$
$=\int_{\Gamma}\int_{\mathcal{X}\times\mathcal{Y}}\mathbb{I}\Big[(\boldsymbol{\hat{\theta}}_{N,\gamma}(x)=\boldsymbol{\theta})\Big]dP^{N}dP^{\gamma},$
@@
Intuitively, the reference distribution captures the full estimation uncertainty by marginalizing over both data uncertainty (variation across datasets $\mathcal{D}_{N}\sim P^{N}$) and procedural uncertainty (variation across $\gamma\sim P^{\gamma}$). Its variance decomposes exactly according to Eqn. 8, enabling direct comparison with the epistemic estimates produced by any second-order method. In practice, the reference distribution is estimated by training models on $n_{d}$ independent datasets with $n_{\gamma}$ random initializations each, yielding $n_{d}\times n_{\gamma}$ models whose empirical variance approximates the theoretical quantity. For real-world datasets where sampling from the data-generating distribution is infeasible, disjoint partitioning provides a practical approximation (see Appendix C.1).
Running Example - Deep Ensembles We evaluate a Deep Ensemble of $10$ members on the synthetic data example described in Eqn. 6. The reference distribution is estimated using $n_{d}=20$ resampling steps and $n_{\gamma}=10$ separate models are trained. The full experimental details are provided in Appendix C.1. Figure 4 illustrates that Deep Ensembles capture only procedural uncertainty, underestimating the variance from the reference distribution.
Figure 4: Comparison of reference distribution and Deep ensemble estimates. The confidence intervals of procedural and data variance as estimated by the reference distribution are shown in blue, while the Deep Ensemble estimate of epistemic uncertainty are shown in green.
Further results on synthetic and real-world data for additional second-order models (Appendix C.2-C.3) reinforce our arguments: Figure 9 demonstrates that across all evaluated methods, epistemic uncertainty estimates fall systematically below the estimate of the reference distribution. Notably, Bayesian methods with richer posteriors produce higher epistemic estimates in low data regimes. However, this rather reflects their failure to learn the underlying data-generating distribution than the correct variance.
As bootstrapped ensembles are trained on resampled versions of $\mathcal{D}_{N}$, one might expect them to better capture the data uncertainty $\text{Var}_{\mathcal{D}_{N}}(\mathbb{E}_{\gamma}[\hat{y}|\boldsymbol{x}]|\mathcal{D}_{N})$, since under regularity conditions, the bootstrap variance consistently approximates the true sampling variance (Efron, 1992). Recent work (Jain and Bates, 2025) decomposes bootstrap-based epistemic uncertainty estimates into data and procedural components. However, bootstrap-based estimates are inherently limited by the empirical support of the dataset and the estimation is only approximate in the finite sample regime. In our experiments (Appendix C.2-C.3) we find that the total epistemic uncertainty as given by the reference distribution is still underestimated.
4 Alternative views
Our position intersects with several lines of critical and constructive work. We review the dominant paradigm of second-order uncertainty representations and its documented limitations. We also examine popular alternative frameworks.